OpenStudy (anonymous):

Researchers have obtained promising results using Cu-67 in the radiotherapy of a number of cancers including non-Hodgkin's lymphoma. If the half-life of this isotope is 2.58 days, by what factor will its activity have decreased after 6.0 days?

OpenStudy (aaronq):

find k, the decay constant: \[t _{1/2}=\frac{ \ln2 }{ k }\] t1/2= half life k=decay constant then find At, \[A _{t}=A _{0}*e ^{-kt}\] At= amount after time elapsed Ao=initial amount k=decay constant (from first equation) t= time elapsed then make a ratio At/Ao to find the factor by which is remains to find he factor by which is decreased subtract that from 1 i.e. 1-At/Ao

OpenStudy (anonymous):

thanks your the best!

OpenStudy (aaronq):

no prob dude